# Hyperons during proto-neutron star deleptonization and the emission of dark flavored particles

### v1

@@ 4 Simulations of the PNS deleptonization with hyperons
For the simulations of the PNS deleptonization, the AGILE-BOLTZTRAN spherically symmetric general relativistic neutrino radiation hydrodynamics model is employed (c.f. Refs. [99, 100, 101, 102] and references therein). It is based on six-species Boltzmann neutrino transport and a complete set of weak interactions (the weak reactions used in the current study, including the references, are listed in Table 1 of Ref. [103]). We distinguish between $\mu$ and $\tau$ (anti)neutrinos due to the inclusion of weak reactions involving muons and antimuons. These include the muonic charged current reactions based on the full kinematics approach [104] as well as neutrino-(anti)muon scattering. For both of which we are following the implementation of Ref. [105].
(a)  18 M
@@
Figure 3:  Post bounce evolution of selected central quantities, comparing simulations with the NUCLEONS (black lines) and HYPERONS (blue lines) EOS showing from left to right, central restmass density $\rho$, central and maximum temperatures, $T_{\rm central}$ and $T_{\rm max}$, central electron and muon abundances, $Y_{e}$ and $Y_{\mu}$, and minimum $Y_{e}$, and the mass fractions $X_{i}$ of neutrons (thick solid lines), protons (thick dashed lines) in comparison to the strange hadrons $\Lambda$ (thin solid lines), $\Sigma^{-}$ (thin dashed lines) and $\Xi$ (thin dash.-dotted lines). Note that the latter are magnified by one order of magnitude for better visibility.
@@
AGILE-BOLTZTRAN has a flexible EOS module that contains the Lattimer & Swesty Skyrme model [106] as well as the comprehensive RMF EOS catalogue of Ref. [107]. The latter incorporates the modified nuclear statistic equilibrium (NSE) with several thousand nuclei [108], including excited state contributions, modeled via temperature dependent statistical weight factors for each nuclear species. The transition from the modified NSE to homogeneous nuclear matter is taken into account via a first-order phase transition construction where the nuclear states are suppressed through a geometric excluded volume approach (for alternative excluded volume functionals, c.f. Ref. [109] and references therein). In Ref. [107], the excluded volume parameter is chosen to result in the complete disappearance of all nuclei at nuclear saturation density. In the present study, we replace the supersaturation density phase of the DD2 RMF EOS [98, 107], with density-dependent meson-nucleon mean-field couplings, with the NUCLEONS and HYPERONS settings from FSU2H*. We find that this approach, despite its ad hoc nature, results in a smooth transition. EOS contributions from electrons, positron and photons are implemented based on the routines of Ref. [110].
@@
The inclusion of the HYPERONS EOS is connected with the implementation of the associated hyperon degrees of freedom, $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$, into the AGILE-BOLTZTRAN CCSN model. It concerns their abundances as well as the single particle properties such as the effective masses, the chemical potentials and the mean-field potentials, according to the underlying model as was discussed in Sec. 3. In particular, the latter will become relevant when implementing weak processes involving hyperons, which, for the present study, are being omitted as the simulation times are limited to $\mathcal{O}(10~{}{\rm s})$ when neutrinos still decouple at lower densities where no hyperons are abundant. This situation will change towards later times, when neutrinos start to decouple in the regions where hyperons are abundant, such that the inclusion of weak processes involving hyperons would influence the further deleptonization and cooling behaviour. Furthermore, the presence of hyperons modifies the charge-neutrality condition, with the negatively charged hyperons considered, i.e. $Y_{e}+Y_{\mu}=Y_{p}-Y_{\Sigma^{-}}-Y_{\Xi^{-}}$. Note that the heavier, neutral hyperons, $\Sigma^{0}$ and $\Xi^{0}$, and the positively charged hyperons, $\Sigma^{+}$ and $\Xi^{+}$, are neglected in the following discussions, as they are all strongly Boltzmann suppressed under the generally extremely neutron-rich conditions at highest densities encountered at the PNS interior, i.e. the following hierarchy applies $Y_{\Lambda}>Y_{\Sigma^{-}}>Y_{\Xi^{-}}$ as well as $Y_{\Sigma^{-}}\gg Y_{\Sigma^{0}}>Y_{\Sigma^{+}}$ and $Y_{\Xi^{-}}\gg Y_{\Xi^{0}}>Y_{\Xi^{+}}$.
@@
With this enhanced EOS module setup, simulations of CCSN are launched from two progenitor models, with zero-age main sequence masses of 18 M and 25 M, both from the stellar evolution series of Ref. [111]. These stellar progenitors are being evolved through all CCSN phases self consistently. It is found that up to several 100 ms post bounce, the abundance of hyperons remains small, below 1%, and no impact could be observed. The simulations with the NUCLEONS and HYPERONS EOS proceed identically.
@@
Table 3: Encloses PNS baryonic and gravitational masses, $M_{\rm B}$, and $M_{\rm G}$, and radius $R$.
progenitor mass1 [[[[${}_{\odot}]$ EOS $M_{\rm B}^{2}$ [[[[${}_{\odot}]$ $M_{\rm G}^{2}$ [[[[${}_{\odot}]$ $R^{2}$ [[[[km]]]]
@@
1
 models from the stellar evolution series of Ref. [111]
2 Evaluated at $\rho=10^{10}$ g cm-3 at about 10 s post bounce
The post-bounce evolution of the four models is illustrated in Fig. 3, for 18 M and 25 M model in Figs. 3(a) and 3(b), respectively, distinguishing the NUCLEONS (black lines) and HYPERONS (blue lines) EOS. Once the supernova explosion proceeds with the shock revival, mass accretion ceases and the enclosed baryon masses change only marginally, which is the ejection of the neutrino driven wind, initially on the order of $10^{-3}$ M s-1 and later substantially lower with $10^{-4}$ M s-1. The resulting PNS masses of the two sets of simulations are listed in Table 3 for both NUCLEONS and HYPERONS runs. These values are evaluated at the restmass density of $10^{11}$ g cm-3 at about 10 s post bounce. The larger PNS masses for the more massive progenitor are due to the higher post-bounce mass accretion rates for the 25 M progenitor models, in particular during the early post-bounce phase, i.e., before the shock revival, as is illustrated in the bottom panel of Fig. 4, showing the mass accretion rate $\dot{M}_{\rm PNS}$, as well evaluated at $\rho=10^{11}$ g cm-3. The top panel illustrates the different PNS radii, $R_{\rm PNS}$, and their evolution for all simulations with the NUCLEONS and HYPERONS EOS under investigation, showing the slightly faster contraction of the HYPERONS PNS due to the softer supersaturation density behaviour.
@@ 2 Evaluated at $\rho=10^{10}$ g cm-3 at about 10 s post bounce
Only after several seconds, in both simulations, differences due to the inclusion of hyperons arise. These differences are best seen by the evolution of the central densities, shown in the left panels of Figs. 3(a) and 3(b), where the simulations based on the HYPERONS EOS reach significantly higher values than the ones with NUCLEONS. This EOS softening is due to the continuous rise of the hyperon abundances for $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$, shown in the right panels. Note that these hyperons are absent for the NUCLEONS simulations. The first hyperons which appear are $\Lambda$’s, due to the lowest mass (see Figs. 1 and 2), already at a density slightly in excess of nuclear saturation density, i.e. already during the early post-bounce evolution prior to the supernova explosion onset. However, the abundance of $\Lambda$ remains low, on the order of $Y_{\Lambda}\simeq 0.001-0.025$ for the for 18 M model and $Y_{\Lambda}\simeq 0.01-0.07$ for the for 25 M model. The latter reach higher abundances because of the generally higher densities and temperatures reached at the central fluid elements in the simulations. Note therefore the different evolution of the central and maximum temperatures in Figs. 3(a) and 3(b). The abundances of the other hyperons, $\Sigma^{-}$ and $\Xi^{-}$, remain about one order of magnitude lower than those of $\Lambda$, in both simulations.
@@
Figure 5:  Post bounce evolution of the neutrino luminosities and average energies for the two sets of simulations launched from 18 M (left panel) and 25 M progenitors (right panel), comparing in addition to the NUCLEONS (black lines) and HYPERONS (blue lines) EOS also those with the inclusion of losses through the emission of dark sector bosons $X^{0}$, denoted as HYPERONS (orange lines). The luminosity in graphs (a) and (b) distinguish the total neutrino luminosity (top panels, solid lines) in comparison to the dark sector particle, denoted as $X^{0}$ (dashed lines), electron flavors (middle panels) and heavy lepton flavors (bottom panel). The average neutrino energies in graphs (c) and (d) show the $\nu_{e}$ (solid lines), $\bar{\nu}_{e}$ (dash dotted lines) and $\nu_{\mu/\tau}$ (dashed lines) as representative for all heavy lepton flavors.
Another difference between the runs with NUCLEONS and HYPERONS is the feedback to the nucleons. As was already discussed in Sec. 3, in particular the abundance of neutrons is reduced due to the appearance of $\Lambda$. This phenomenon is observed consistently on the CCSN simulations too. The neutron abundances are shown in the right panels of Figs. 3(a) and 3(b). The reduced neutron abundance in addition softens the high-density EOS, which is governed by neutrons with a dominating abundance of $Y_{n}\simeq 0.75-0.7$. As a consequence of the different hadronic abundances, in order to obey the condition of charge neutrality, there is a feedback to the electron and muon abundances. In particular, the central muon abundance is reduced substantially for the massive 25 M model. Note also that the central muon abundance is substantially lower and that the central electron abundance is substantially higher than in the EOS discussion in Sec. 3, where $\beta$-equilibrium was assumed. The observed difference of the electron and in particular the muon abundances are a consequence of the lower central temperatures obtained for the simulations based on the HYPERONS EOS (see Fig. 3), since the muons are produced thermally through the muonic charged current processes.
@@
Despite the differences obtained at the PNS interior, structure and stability of the PNS are not affected by the presence of hyperons. This is partly related to the conditions for the appearance of hyperons, which in turn gives rise to finite hyperon abundances, in particular $\Lambda$ only at the very central fluid elements in the simulations. As a consequence, the locations of the neutrinospheres are not affected by the presence of hyperons and hence the neutrino luminosities and average energies are indistinguishable between the two runs NUCLEONS and HYPERONS, for both intermediate-mass 18 M model, shown in Figs. 5(a) and 5(c), and high-mass 25 M model, shown in Figs. 5(b) and 5(d). The evolution of the neutrino luminosities and average energies shown in Fig. 5 are sampled in the co-moving frame of reference at a distance of 500 km.
5 Dark flavored sectors during PNS deleptonization
@@
The results including the dark flavored sector are based on the same HYPERON EOS and will be henceforth denoted as HYPERONS, which will be discussed in this section. To investigate the impact of a dark flavored sector in CCSN, we need to estimate the emissivity in the new channels. We assume that the emitted particles are massless (i.e. their mass is much smaller than $T_{\rm PNS}$) and that the interactions are such that their mean free path is much larger than the radius of the PNS. These particles will then freely stream out of the PNS once they are produced in its core.
5.1 Dark flavored emissivity
@@
For concreteness, in the following we will focus on the emission of neutral dark bosons $X^{0}$, such as an axion or dark photon, produced by the flavor-violating transition $s\to d+X^{0}$. This is parametrized by the following Lagrangians,
@@
where $a$ is the axion field and $f_{a}$ its decay constant and with strange and down quark spinors denoted as ${\rm s}$ and ${\bar{\rm d}}$, respectively. $F_{\mu\nu}^{\prime}$ is the field strength tensor of the dark photon and $\Lambda$ the energy scale associated to the UV completion of the effective Lagrangian $\mathcal{L}_{\gamma^{\prime}}$, and $c^{V,A}_{\rm ds}$ and $c^{T,T5}_{\rm ds}$ are dimensionless (effective) couplings. The main hadronic production mechanism in the PNS induced by Lagrangians (5.1) and (5.2) will then be $\Lambda\to n+X^{0}$, whereby decays involving heavier hyperons, such as $\Sigma^{+}\to p+X^{0}$, are suppressed by their lower abundances and neglected in the present analysis. Following Ref. [72], the spectrum of the energy loss rate of the nuclear medium per unit volume, denoted as $Q$, with respect to the $X_{0}$ energy $E_{a}$, in the PNS rest frame and induced by the process $\Lambda\to n+X^{0}$ is given by the following expression,
@@
with ${\bar{E}}=(m_{\Lambda}^{2}-m_{n}^{2})/2m_{\Lambda}$, $E$ is the energy of the $\Lambda$ with $E_{0}=m_{\Lambda}(E_{a}^{2}+{\bar{E}}^{2})/(2E_{a}{\bar{E}})$ and $f_{\Lambda,n}$ are the relativistic Fermi-Dirac phase-space distribution functions of $\Lambda$ and $n$. The information of the transition amplitude is encoded in $\Gamma_{X}$, which is the decay rate of $\Lambda\to n+X^{0}$ in vacuum,
$\Gamma_{X}=\frac{\bar{E}_{a}^{3}C_{X}}{2\pi}~{},$ (5.4)
with $C_{a}=(f_{1}^{2}|c_{\rm ds}^{V}|^{2}+g_{1}^{2}|c_{\rm ds}^{A}|^{2})/4f_{a}^{2}$, where $f_{1}=-1.22(6)$ and $g_{1}=-0.89(2)$ are the vector and axial-vector baryonic couplings determined in [72], and $C_{\gamma^{\prime}}=8g_{T}^{2}(|c_{\rm ds}^{T}|^{2}+|c_{\rm ds}^{T5}|^{2})/\Lambda^{2}$, where $g_{T}\approx-0.73$ is the baryonic tensor coupling [further details can be found in 73, and references therein].
@@
$\epsilon\approx\delta\;X_{\Lambda}\tau_{\Lambda}^{-1}\,{\rm BR}(\Lambda\to nX^{0})~{},$ (5.5)
where $X_{\Lambda}$ is the abundance of $\Lambda$ and where we have re-expressed the decay width in terms of the lifetime, $\tau_{\Lambda}$, and the branching fraction of the decay, BR. Adopting the classical upper limit of the emissivity, $\epsilon_{\rm max}=10^{19}$ erg s-1 g-1, from SN1987A at the conditions of the PNS predicted at about 1 second post-bounce, one obtains,
$\epsilon\approx\epsilon_{\rm max}\left(\frac{X_{\Lambda}}{0.01}\right)\left(\frac{1.6\times 10^{-9}}{\text{BR}(\Lambda\to nX^{0})}\right)~{}.$ (5.6)
This leads, for a characteristic $\Lambda$ abundance of 1%, to an upper limit of the branching fraction in the range of $10^{-9}$, which overestimates the conservative limit by a factor $\approx 5$, which was obtained using the full expression in Eq. (5.3) and state-of-the-art spherically symmetric simulations reported in Ref. [64]. Nevertheless, this upper bound is many orders of magnitude stronger than those that can be obtained from laboratory experiments, and, as was found in Refs. [72, 73], it leads to the strongest constraint on the axial couplings of the flavored QCD axion and on the massless dark photon couplings. In addition, as discussed in Ref. [73], in the case of $\Lambda$ processes one would still produce a too large emission in the deep trapping regime as the dark sector luminosity would stem from the last surface where $\Lambda$’s can coexist in equilibrium in the plasma and which corresponds to a region of high temperatures.
5.2 Impact of dark flavored sector on PNS evolution
The non-relativistic emissivity (5.5) is implemented into the AGILE-BOLTZTRAN CCSN model as sink for the internal energy equation, following the implementation of dark boson losses of Refs. [113, 57]. Here, we use the literature branching ratios of $10^{-8}$ for the intermediate-mass 18 M model and $10^{-9}$ for the massive 25 M model, in accordance with Ref. [73]. It results in enhanced cooling contributions in the domain where $\Lambda$ are abundant. This is the case only at the innermost 15–20 km of the PNS, as is illustrated for both 18 M and 25 M models in Fig. 6 for two distinct post-bounce times of 1 s and 5 s, comparing the three setups NUCLEONS (black lines), HYPERONS (blue lines) and HYPERONS (orange lines).

### v2

@@ 4 Simulations of the PNS deleptonization with hyperons
For the simulations of the PNS deleptonization, the AGILE-BOLTZTRAN spherically symmetric general relativistic neutrino radiation hydrodynamics model is employed (c.f. Refs. [112, 113, 114, 115] and references therein). It is based on six-species Boltzmann neutrino transport and a complete set of weak interactions (the weak reactions used in the current study, including the references, are listed in Table 1 of ref. [116]). We distinguish between $\mu$ and $\tau$ (anti)neutrinos due to the inclusion of weak reactions involving muons and antimuons. These include the muonic charged current reactions, based on the full kinematics approach [117], as well as neutrino-(anti)muon inelastic scattering. For both of which we are following the implementation of ref. [118].
(a)  18 M progenitor
@@
Figure 3:  Post bounce evolution of selected central quantities for the 18 M and 25 M models in graph (a) and (b) respevtively, comparing simulations with the NUCLEONS (black lines) and HYPERONS (blue lines) EOS showing from left to right, central restmass density $\rho$, central and maximum temperatures, $T_{\rm central}$ and $T_{\rm max}$, central electron and muon abundances, $Y_{e}$ and $Y_{\mu}$, and minimum $Y_{e}$, and the mass fractions $X_{i}$ of neutrons in the top panel (thick solid lines) and in the bottom panel protons (thick dotted lines) in comparison to the strange hadrons $\Lambda$ (thin solid lines), $\Sigma^{-}$ (thin dashed lines) and $\Xi$ (thin dash-dotted lines). Note that the latter are magnified by one order of magnitude for better visibility.
@@
AGILE-BOLTZTRAN has a flexible EOS module that contains the Lattimer & Swesty Skyrme model [119] as well as the comprehensive RMF EOS catalogue of ref. [120]. The latter incorporates the modified nuclear statistic equilibrium (NSE) with several thousand nuclei [121], including excited state contributions, modelled via temperature dependent statistical weight factors for each nuclear species. The transition from the modified NSE to homogeneous nuclear matter is taken into account via a first-order phase transition construction where the nuclear states are suppressed through a geometric excluded volume approach (for alternative excluded volume functionals, c.f. ref. [122] and references therein). In ref. [120], the excluded volume parameter is chosen to result in the complete disappearance of all nuclei at nuclear saturation density. In the present study, we replace the supersaturation density phase of the DD2 RMF EOS [111, 120], with density-dependent meson-nucleon mean-field couplings, with the NUCLEONS and HYPERONS settings from FSU2H. We find that this approach, despite its ad hoc nature, results in a smooth transition. EOS contributions from electrons, positron and photons, as well as Coulomb contributions for the non-NSE regime, are implemented based on the routines of ref. [123].
@@
The inclusion of the HYPERONS EOS is connected with the implementation of the associated hyperon degrees of freedom, $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$, into the AGILE-BOLTZTRAN CCSN model. It concerns their abundances as well as the single particle properties such as the effective masses, the chemical potentials and the mean-field potentials, according to the underlying model as was discussed in Sec. 3. In particular, the latter will become relevant when implementing weak processes involving hyperons, which, for the present study, are being omitted as the simulation times are limited to $\mathcal{O}(10~{}{\rm s})$ when neutrinos still decouple at lower densities where no hyperons are abundant. This situation will change towards later times, when neutrinos start to decouple in the regions where hyperons are abundant, such that the inclusion of weak processes involving hyperons would influence the further deleptonization and cooling behaviour. Furthermore, the presence of hyperons modifies the charge-neutrality condition, with the negatively charged hyperons considered, i.e. $Y_{e}+Y_{\mu}=Y_{p}-Y_{\Sigma^{-}}-Y_{\Xi^{-}}$. Note that despite their consistent treatment and implementation into the CCSN simulations, the heavier, neutral hyperons, $\Sigma^{0}$ and $\Xi^{0}$, and the positively charged hyperons, $\Sigma^{+}$ and $\Xi^{+}$, are neglected in the following discussions, as they are all strongly Boltzmann suppressed under the extremely neutron-rich conditions at highest densities encountered at the PNS interior, i.e. the following abundance hierarchy generally holds: $Y_{\Lambda}>Y_{\Sigma^{-}}>Y_{\Xi^{-}}$ as well as $Y_{\Sigma^{-}}\gg Y_{\Sigma^{0}}>Y_{\Sigma^{+}}$ and $Y_{\Xi^{-}}\gg Y_{\Xi^{0}}>Y_{\Xi^{+}}$.
@@
With this enhanced EOS module setup, simulations of CCSN are launched from two progenitor models, with zero-age main sequence masses of 18 M and 25 M, both from the stellar evolution series of ref. [124]. These stellar progenitors are being evolved through all CCSN phases self consistently. It is found that up to several 100 ms post bounce, the abundance of hyperons remains small, below 1%, and no impact could be observed, not on the overall dynamics nor on the neutrino heating and cooling rates. The neutrino signals are indistinguishable between NUCLEONS and HYPERONS.
The simulations with the NUCLEONS and HYPERONS EOS proceed identically.
@@
Table 4: Stellar progenitor model’s zero-age main sequence masses, $M_{\rm ZAMS}$, and the encloses PNS baryonic and gravitational masses, $M_{\rm B}$, and $M_{\rm G}$, as well as the corresponding radii $R$, evaluated at the end of the CCSN simulations.
progenitor massa EOS $M_{\rm B}^{\rm b}$ $M_{\rm G}^{\rm b}$ $R^{\rm b}$
$M_{\rm ZAMS}\,[$${}_{\odot}]$ [[[[${}_{\odot}]$ [[[[${}_{\odot}]$ [[[[km]]]]
@@
a models from the stellar evolution series of ref. [124]
b Evaluated at $\rho=10^{10}$ g cm-3 at about 10 s post bounce
The post-bounce evolution of the four models is illustrated in Fig. 3, for 18 M and 25 M model in Figs. 3(a) and 3(b), respectively, distinguishing the NUCLEONS (black lines) and HYPERONS (blue lines) EOS. Once the supernova explosion proceeds with the revival of the shock, mass accretion onto the PNS surface ceases and the enclosed PNS baryon masses change only marginally. The latter is due to the ejection of the neutrino driven wind, initially on the order of $10^{-3}$ M s-1 and later substantially lower with $10^{-4}$ M s-1. The resulting PNS masses of the two sets of simulations are listed in Table 4, for both NUCLEONS and HYPERONS runs. These values are evaluated at the restmass density of $10^{11}$ g cm-3 at about 10 s post bounce. The larger PNS masses for the more massive progenitor are due to the higher post-bounce mass accretion rates for the 25 M models, in particular during the early post-bounce phase, i.e., before the shock revival, as is illustrated in the bottom panel of Fig. 4, showing the mass accretion rate $\dot{M}_{\rm PNS}$, as well evaluated at $\rho=10^{11}$ g cm-3. The top panel illustrates the different PNS radii, $R_{\rm PNS}$, and their evolution for all simulations with the NUCLEONS and HYPERONS EOS under investigation, showing the slightly faster contraction of the HYPERONS PNS due to the softer supersaturation density behaviour.
@@
Only after several seconds, in both simulations, differences due to the inclusion of hyperons arise. These differences are best seen by the evolution of the central densities, shown in the left panels of Figs. 3(a) and 3(b), where the simulations based on the HYPERONS EOS reach significantly higher values than the ones with NUCLEONS. This EOS softening is due to the continuous rise of the hyperon abundances for $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$, shown in the right panels. Note that these hyperons are absent for the NUCLEONS simulations (black lines). The first hyperons which appear are $\Lambda$’s, due to the lowest mass (see Figs. 1 and 2), already at a density slightly in excess of nuclear saturation density, i.e. already during the early post-bounce evolution prior to the supernova explosion onset. However, the abundance of $\Lambda$ remains low, on the order of $Y_{\Lambda}\simeq 0.001-0.025$ for the for 18 M model and $Y_{\Lambda}\simeq 0.01-0.07$ for the for 25 M model. The latter reach higher abundances because of the generally higher densities and temperatures obtained at the central fluid elements in the simulations. Note therefore the different evolution of the central and maximum temperatures in Figs. 3(a) and 3(b). The abundances of the other hyperons, $\Sigma^{-}$ and $\Xi^{-}$, remain about one order of magnitude lower than those of $\Lambda$, in both simulations.
@@
Figure 5:  Post bounce evolution of the neutrino luminosities and average energies for the two sets of simulations launched from 18 M (left panel) and 25 M progenitors (right panel), comparing in addition to the NUCLEONS (black lines) and HYPERONS (blue lines) EOS also those with the inclusion of losses through the emission of dark sector bosons $X_{0}$, denoted as HYPERONS (orange lines). The luminosity in graphs (a) and (b) distinguish the total neutrino luminosity (top panels, solid lines) in comparison to the dark sector particle, denoted as $X_{0}$ (dashed lines), electron flavours (middle panels) and heavy lepton flavours (bottom panel). The average neutrino energies in graphs (c) and (d) show the $\nu_{e}$ (solid lines), $\bar{\nu}_{e}$ (dash dotted lines) and $\nu_{\mu/\tau}$ (dashed lines) as representative for all heavy lepton flavours.
[…]
@@
Despite the differences obtained at the PNS interior, structure and stability of the PNS are not affected by the presence of hyperons. This is partly related to the conditions for the appearance of hyperons, which in turn gives rise to finite hyperon abundances, in particular $\Lambda$ only at the very central fluid elements in the simulations. As a consequence, the locations of the neutrinospheres are not affected by the presence of hyperons and hence the neutrino luminosities and average energies are indistinguishable between the runs NUCLEONS and HYPERONS, for the intermediate-mass 18 M model, shown in Figs. 5(a) and 5(c), and high-mass 25 M model, shown in Figs. 5(b) and 5(d). The evolution of the neutrino luminosities and average energies shown in Fig. 5 are sampled in the co-moving frame of reference at a distance of 500 km.
5 Dark flavoured sectors during PNS deleptonization
@@
The results including the dark flavoured sector are based on the same HYPERON EOS and will be henceforth denoted as HYPERONS, which will be discussed in this section. To investigate the impact of a dark flavoured sector in CCSN, we need to estimate the emissivity in the new channels. We assume that the emitted particles are massless (i.e. their mass is much smaller than $T_{\rm PNS}$) and that the interactions are such that their mean free path is much larger than the radius of the PNS. These particles will then freely stream out of the PNS once they are produced in its core.
5.1 Dark flavoured emissivity
@@
For concreteness, in the following we will focus on the emission of neutral dark bosons $X_{0}$, such as an axion or dark photon, produced by the flavour-violating transition $s\to d+X_{0}$. This is parametrised by the following Lagrangians,
@@
where $a$ is the axion field and $f_{a}$ its decay constant and with strange and down quark fields denoted as ${\rm s}$ and ${\bar{\rm d}}$, respectively. $F_{\mu\nu}^{\prime}$ is the field strength tensor of the dark photon and $\Lambda$ the energy scale associated to the UV completion of the effective Lagrangian $\mathcal{L}_{\gamma^{\prime}}$ (see ref. [97] for effective vectorial couplings of the dark photon and some UV completions), and $c^{V,A}_{\rm ds}$ and $c^{T,T5}_{\rm ds}$ are dimensionless (effective) couplings. The main hadronic production mechanism in the PNS induced by Lagrangians (5.1) and (5.2) will then be $\Lambda\to n+X_{0}$, whereby decays involving heavier hyperons, such as $\Sigma^{+}\to p+X_{0}$, are suppressed by their lower abundances and neglected in the present analysis. Following ref. [82], the spectrum of the energy loss rate of the nuclear medium per unit volume, denoted as $Q$, with respect to the $X_{0}$ energy $E_{a}$, in the PNS rest frame and induced by the process $\Lambda\to n+X_{0}$ is given by the following expression,
@@
with ${\bar{E}}=(m_{\Lambda}^{2}-m_{n}^{2})/2m_{\Lambda}$, $E$ is the energy of the $\Lambda$ with $E_{0}=m_{\Lambda}(E_{a}^{2}+{\bar{E}}^{2})/(2E_{a}{\bar{E}})$ and $f_{\Lambda,n}$ are the relativistic Fermi-Dirac phase-space distribution functions of $\Lambda$ and $n$. The information of the transition amplitude is encoded in $\Gamma_{X}$, which is the decay rate of $\Lambda\to n+X_{0}$ in vacuum,
$\Gamma_{X}=\frac{\bar{E}^{3}C_{X}}{2\pi}~{},$ (5.4)
with $C_{a}=(f_{1}^{2}|c_{\rm ds}^{V}|^{2}+g_{1}^{2}|c_{\rm ds}^{A}|^{2})/4f_{a}^{2}$, where $f_{1}=-1.22(6)$ and $g_{1}=-0.89(2)$ are the vector and axial-vector baryonic couplings determined in ref. [82], and $C_{\gamma^{\prime}}=8g_{T}^{2}(|c_{\rm ds}^{T}|^{2}+|c_{\rm ds}^{T5}|^{2})/\Lambda^{2}$, where $g_{T}\approx-0.73$ is the baryonic tensor coupling (further details can be found in ref. [83] and references therein).
@@
$\epsilon\approx\delta\;Y_{\Lambda}\tau_{\Lambda}^{-1}\,{\rm BR}(\Lambda\to n+X_{0})~{},$ (5.5)
where we have re-expressed the decay width in terms of the lifetime, $\tau_{\Lambda}$, and the branching fraction of the decay, BR. Adopting the classical upper limit of the emissivity, $\epsilon_{\rm max}=10^{19}$ erg s-1 g-1, from SN1987A at the conditions of the PNS predicted at about 1 second post-bounce, one obtains,
$\epsilon\approx\epsilon_{\rm max}\left(\frac{Y_{\Lambda}}{0.01}\right)\left(\frac{1.6\times 10^{-9}}{\text{BR}(\Lambda\to n+X_{0})}\right)~{}.$ (5.6)
This leads, for a characteristic $\Lambda$ abundance of 1%, to an upper limit of the branching fraction in the range of $10^{-9}$, which overestimates the conservative limit by a factor $\approx 5$, which was obtained using the full expression in Eq. (5.3) and state-of-the-art spherically symmetric simulations reported in ref. [74]. Nevertheless, this upper bound is many orders of magnitude stronger than those that can be obtained from laboratory experiments, and, as was found in Refs. [82, 83], it leads to the strongest constraint on the axial couplings of the flavoured QCD axion and on the massless dark photon couplings. In addition, as discussed in ref. [83], in the case of $\Lambda$ processes one would still produce a too large emission in the deep trapping regime as the dark sector luminosity would stem from the last surface where $\Lambda$’s can coexist in equilibrium in the plasma and which corresponds to a region of high temperatures.
5.2 Impact of dark flavoured sector on PNS evolution
The non-relativistic emissivity (5.5) is implemented into the AGILE-BOLTZTRAN CCSN model as sink for the internal energy equation, following the implementation of dark boson losses of Refs. [126, 67]. Here, we use the literature branching ratios of $10^{-8}$ for the intermediate-mass 18 M model and $10^{-9}$ for the massive 25 M model, in accordance with ref. [83]. It results in enhanced cooling contributions in the domain where $\Lambda$ are abundant 333 These limits can be readily transformed in stringent limits on the flavoured QCD axion or massless dark photon described by the Lagrangians in (5.1) and (5.2). For example, in the QCD axion case with $\mathcal{O}(1)$ off-diagonal flavour couplings, the upper limit $\text{BR}(\Lambda\to n+a)\leq 10^{8}$ corresponds to the limit $f_{a}\gtrsim 5\times 10^{9}\text{ GeV}$ or $m_{a}\lesssim 1\text{ meV}$ (see Refs. [83] and [82] for details). . This is the case only at the innermost 15–20 km of the PNS, as is illustrated for both 18 M and 25 M models in Fig. 6 for two distinct post-bounce times of 1 s and 5 s, comparing the three setups NUCLEONS (black lines), HYPERONS (blue lines) and HYPERONS (orange lines).